OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (anonymous):

you first start off with the formula y-y1=m(x-x1) then you plug in those numbers so it would be y+4=(-3/8)(x-5) then you do distributive property for the slope and x-5 then subtract 4 for standard form!

OpenStudy (anonymous):

So what is it?

OpenStudy (anonymous):

can you figure it out yourself?

OpenStudy (anonymous):

8y+3x=-17 That's what I got.

OpenStudy (anonymous):

you solve the equation for y

OpenStudy (anonymous):

it would be y=-3/8x-17/8

OpenStudy (anonymous):

That doesn't sound right.

OpenStudy (anonymous):

I think *in standard form*, it's 8y+3x=-17

OpenStudy (anonymous):

yeah your right i was thinking of a different way